OpenStudy (anonymous):

Calculus 3: Absolute Maximum & Minimum f(x,y)=2+2x+2y-x^2-y^2 with a closed triangular region with vertices at (0,0), (0,6), and (6,0) I have found the critical point of the interior of the triangle but to find the max and min of the boundaries I just plugin (0,0), (0,6), and (6,0) into f(x,y) to figure it out correct? If so, can the same be said for a rectangle?

OpenStudy (anonymous):

Also, is the absolute maximum f(1,1)=4 and absolute global minimum f(0,9) and f(9,0)?

OpenStudy (amistre64):

you need to determine your partials right?

OpenStudy (amistre64):

f(x,y)=2+2x+2y-x^2-y^2 Fx = 2 - 2x = 0, x = 1 Fxx = -2 Fxy = 0 Fy = 2 - 2y = 0; y = 1 Fyy = -2 D > 0 .... but i tend to forget what that tells us about a possible saddle point :/

OpenStudy (amistre64):

If D>0 and Fxx >0 then there is a relative minimum If D>0 and Fxx < 0 then there is a relative maximum so i believe that tells us that 1,1 is a relative max, is that within or on our boundary region?

OpenStudy (amistre64):

you have to define the line between the vertexes, and evaluate the function along that line ...

OpenStudy (amistre64):

(0,0) (0,6) x=0, y = 0 to 6 f(0,y)=2+2y-y^2 Fy = 2-2y= 0 at y=1 again, so the point (0,3) is a critical point to test ... if im remembering it correctlyt

OpenStudy (anonymous):

So you have to check along the lines and vertices correct?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Ohhhhhh! Seems like more work but definitely doable now =)

OpenStudy (anonymous):

One moment, I'm trying to trying a PDF that found the maximum and minimum but the professor didn't bother with the lines and only focus on the vertices.

OpenStudy (amistre64):